# Feasible Space Monitoring for Multiple Control Barrier Functions with application to Large Scale Indoor Navigation

### v1

@@ V-A2 Chebyshev Ball of a Polytope
The center $c$ and radius $r$ of the largest sphere contained inside a polytope, called Chebyshev ball, can be computed with the following Linear Program (LP)[21, Chapter 8]
@@
V-A3 Inscribing Ellipsoid
An ellipoid $E$ is an image of the unit ball under affine transformation $(B,d)$, $E=\{Bu+d~{}|~{}||u||\leq 1\}$ where $B\in S_{++}^{m}$ and $d\in\mathbb{R}^{m\times 1}$. Using the fact that the ellipsoid volume is proportional to $\det B$, the maximum volume ellisoid can be obtained by solving the following convex optimization problem [21, Chapter 8]
$\min_{B,d}\quad$ $\log\det B$ (9a)
@@ V-A3 Inscribing Ellipsoid
V-B Polytope Volume Gradients
Gradients over Monte Carlo estimate are prone to be noisy to the choice of random samples $s_{i}$. The sphere and ellipsoid on the other hand are computed using convex programs and, therefore, we can exploit sensitivity analysis [22] to compute the gradients of their volume w.r.t the polytope matrices $A,b$. Libraries like cvxpy [23] readily provide interface to differentiate through convex programs. However, we note that the volume of the polytope as well as its inscribing sphere and ellipsoid are non-differentiable quantities. Consider the scenario shown in Fig. 4. Suppose the hyperplanes H1-H4 are stationary and the H5 is free to move towards the origin, thereby decreasing enclosed volume, or away from the origin, having no effect on the enclosed volume. The point of non-differentiability for the circle or ellipsoid is when H6 is tangent to them. For the remainder of the paper, we make the following assumption and leave the non-smooth analysis to future work.
Assumption 1.
The volume of the convex polytope (7) is differentiable w.r.t to its constituent matrices $A,b$.
Remark 1.
In practice, we may encounter states at which the volume is non-differentiable w.r.t matrices $A,b$. Cvxpy returns heuristic gradient values at points of non-differentiability and we employ these in our implementations. In experiments, we did not observe any negative effects of this approximation, but a more formal non-smooth control theoretic analysis and consideration of generalized gradients, similar to non-smooth CBFs[2, 24] is left for future work.
V-C Polytope Volume barrier function
Let $\mathcal{V}(t,x)=\textrm{vol}(\mathcal{U}_{c})(t,x)$. The compatible state space is defined as the set of states $D(t)=\{x\in\mathcal{X}~{}|~{}\mathcal{V}(t,x)\geq 0\}$ and is visualized in Fig. 3. We then aim to find a CBF $\mathcal{V}_{c}:\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$, with class-$\mathcal{K}$ function $\alpha_{\mathcal{V}_{c}}$, that renders $D(t)$ forward invariant using, for example, the following controller
$\pi(t,x)=\arg\min_{u\in\mathcal{U}}\quad$ $||u-u_{ref}(t,x)||^{2}$
s.t. $\dot{\mathcal{V}_{c}}(t,x,u)\geq-\alpha_{\mathcal{V}_{c}}(\mathcal{V}(t,x))$ (10)
Assumption 2.
(Necessity of CBFs) For any controlled invariant set $C\subset\mathcal{X}$ of the (1), there exists a continuously differentiable function $h:\mathbb{R}^{+}\times\mathcal{X}\rightarrow\mathbb{R}$ such that $\mathcal{C}$ is a 0-superlevel set of $h$ and $h$ is a CBF in sense of (1).
Under Assumption 2, the following theorem establishes the necessity and sufficiency of FS-CBF.
@@ V-C Polytope Volume barrier function
Let $h_{i},i\in\{1,..,N\}$ be $N$ barrier functions corresponding to $N$ safe sets defined as in (III). Let $\mathcal{V}(t,x)$ be the volume of the feasible CBF space $\mathcal{U}_{c}(t,x)$ in the control domain defined in (6). Let $D(t)=\{x\in\mathcal{X}~{}|~{}\mathcal{V}(t,x)\geq 0\}$ be the compatible state space of $h_{i},i\in\{1,..,N\}$. Then under Assumption 2, $h_{i}$ are persistently feasible iff there exists $\mathcal{C}\subset\mathcal{D}$ and a function $\mathcal{V}_{c}:\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that $\mathcal{V}_{c}$ is a control barrier function on +×$\mathbb{R}^{+}\times\mathcal{C}$ and $x(0)\in\mathcal{C}$.
@@
We prove the sufficiency first. Suppose that $\mathcal{V}_{c}$ is a CBF with class-$\mathcal{K}$ function $\alpha_{\mathcal{V}}$ on +×$\mathbb{R}^{+}\times\mathcal{C}$, with $\mathcal{C}\subset\mathcal{D}$. If $x(0)\in\mathcal{C}$, then $x(t)\in\mathcal{C},\forall t>0$ under the action of the controller (10). Therefore, $\mathcal{V}_{c}(t,x(t))\geq 0,\forall t>0$ and hence $\mathcal{V}(t,x(t))>0$ for all $t>0$, implying persistent feasibility.
Next, we show the necessity. Suppose $h_{i},i\in\{1,2,..,N\}$ are persistently feasible. This implies that there exists a subset $E\subset D$ that is control invariant. Existence of $\mathcal{V}_{c}$ now trivially follows from Assumption 2. ∎
Example 1.
We demonstrate a procedure to find a valid FS-CBF for navigating in the obstacle environment shown in Fig. 2. A Dubins car model is considered: $\dot{x}=v\cos\theta,\dot{y}=v\sin\theta,\dot{\theta}=u$ where $x,y$ are position, $\theta$ is heading, $v$ is fixed speed, and the angular velocity $u\in[-0.5,0.5]$ is the control input. We consider the constraint functions $h_{i}=d_{i}^{2}-d_{min}^{2}$, with $d_{i}$ being the distance to center of $i^{th}$ obstacle and $d_{min}$ the radius of obstacle. The compatible state space is thus given as the states $x,y,\theta$ for which the feasible space (defined by the control input bounds and the corresponding Higher-Order CBF (HOCBF) constraints) is non-zero. $\mathcal{U}_{c}(x)$ is visualized in Fig.5 for two different choices of linear class-$\mathcal{K}$ functions in the HOCBFs. To find the CBF corresponding to largest invariant set inside the compatible state space in Fig. 3(c), we use the refineCBF framework from [25] that leverages Hamilton-Jacobi (HJ) reachability analysis. A grid of $201\times 201\times 81$ is used to discretize the state space in the domain $x,y\in[0,4],\theta\in[-\pi,\pi]$, and a time horizon of 10 for computing the backward reachable sets. The feasible space volume is set to $-1000$ whenever $\min(h_{1},h_{2},h_{3})<0$. The constructed FS-CBF space is shown in Fig. 5(b),(d). It is observed that the compatible state space is shrunk to remove states from where compatibility cannot be ensured in future with any controller.
Figure 5: Compatible space vs refined FS-CBF values at different $X,Y$ and heading $\theta=45^{0}$. (a),(c) Feasible space volume values for Case 1 - aggressive (a) and Case 2 - conservative (b) class-$\mathcal{K}$ function for HOCBF. (b), (d) refined CBF values. The compatible space (purple) is a subset of safe CBF space (yellow).
Example 1 shows an application of an existing CBF synthesis method to construct $\mathcal{V}_{c}$ for a simple system. For highly dynamic and constrained environments though, as previously mentioned, it is usually computationally prohibitive to consider all possible scenarios in the design of CBFs. Therefore, in practice, we assume $\mathcal{V}_{c}=\mathcal{V}$ and use the controller (10) to enforce persistent feasibility.
Remark 2.
We will demonstrate later in Section VI-A and Fig. 6 that imposing constraint (10) results in a control being chosen in the interior of feasible space of constraints (10), even when the nominal control input is outside the feasible space. This is done to prevent the volume from shrinking too fast. The controller (4) on the hand always chooses a point on boundary of feasible space. A mapping from $u_{ref}$ to the interior of feasible space was also designed in [26] through a gauge map. The gauge map was designed in a different way than ours and focused on obtaining differentiable control design with a single barrier constraint only.
V-D Valid class-$\mathcal{K}$ function for Feasible Space CBF
Finding a valid CBF entails simultaneously finding a barrier function $h$ and a class-$\mathcal{K}$ function $\alpha$ that satisfies the CBF derivative condition. While linear class-$\mathcal{K}$ functions are the first choice, below we show through two examples that the structure of the manifold defined by feasible space enclosed by a polytope can preclude the existence of such a function and propose a class-$\mathcal{K}$ function that it does admit. This analysis also provides insight into the nature of CBF-based controllers.
Example 2.
Consider a symmetric polygon of side length $r$ in $m$-dimensional space, visualized in Fig. 3(d) for the 2D case. The volume of the polygon (or of its inscribing sphere) can easily be found to be $kr^{m}$, where $k>0$ depends on the dimension $m$. Suppose all the constituent hyperplanes move at a rate $\dot{r}_{v}$ leading to a shrinking enclosed volume. Then, using $r=(\mathcal{V}/k)^{1/m}$, we have
$\dot{\mathcal{V}}=kmr^{m-1}\dot{r}=km(k\mathcal{V})^{m-1/m}\dot{r}_{v}$ (11)
where $\dot{r}_{v}<0$. For the volume to be a CBF, we require a function $\alpha$ such that
$km\dot{r}(k\mathcal{V})^{m-1/m}\geq-\alpha(\mathcal{V})\implies km|\dot{r}|(k\mathcal{V})^{m-1/m}\leq\alpha(\mathcal{V})$
A linear class-$\mathcal{K}$ function cannot satisfy the above condition. A candidate function is $\alpha(z)\propto(\mathcal{V})^{m-1/m}$.
@@ VI Simulation Results
In this section, we present two case studies to evaluate the performance of the proposed volume barrier function.
VI-A Case Study 1: Improving Feasibility and Sensitivity
@@
Figure 6: Case study 1 - humans move towards the right. (a),(b) Paths for the three cases of $k_{v}=1,1.5,2.5$. Blue and red paths in (a) end when CBF-QP becomes infeasible. (c), (d), (e), (f) Feasible space visualization with and without volume barrier function for $k_{v}=1.5$ in (c), (d) and $k_{v}=2.5$ in (e), (f). Note when nominal input is outside the feasible space, the volume barrier function maps it to interior of feasible space whereas CBF-QP (4) maps it to a point on the boundary.
@@
This case study serves two purposes. First, we show that the addition of a feasible space volume barrier function enhances persistent feasibility. Second, it also helps reduce sensitivity to the hyperparameters of the controller. Consider the scenario shown in Fig. 6 for a robot navigating to its goal location $g$ while avoiding black prohibited zones and performing collision avoidance with humans. The robot is modeled as a dynamic unicycle with position $p=[p_{x},p_{y}]$, velocity $v$, heading angle $\psi$ and the following dynamics $\dot{p}_{x}=v\cos\psi,~{}\dot{p}_{y}=v\sin\psi,~{}\dot{v}=a,~{}\dot{\psi}=\omega$. The control inputs are the linear acceleration $a$ and angular velocity $\omega$. The humans move with social force model [27][28], and consider the robot as one of the social agents. We consider the following reference (nominal) controller for the extended unicycle dynamics
@@
where $k_{\omega},k_{x},k_{v}>0$ are user-chosen gains. We implement the CBF-QP (4) for three different values of the gain $k_{v}$ to do sensitivity analysis. Cvxpylayers[23] is used to solve for the ellipse volume, and compute its gradient w.r.t polytope equations matrices $A,b$, and JAX’s autograd feature [29] is used to compute the gradient of $A,b$ w.r.t states of the robot, humans, and the obstacles. We use JAX’s JIT feature to speed up computations whenever possible, and each control input computation takes 0.03 seconds. Fig. 6(a)-(b) show the trajectories of the robot, and Figs. 6(c)-6(e) show the variation of polytope volume under three different values of the gain $k_{v}$. We observe that when $k_{v}$ is increased, that is, the nominal controller is more aggressive, the CBF-QP becomes infeasible sooner. Now, we impose the constraint that restricts the rate-of-change of the ellipse volume (10), computed using (V-A3) as a CBF condition. Figs. 6(d), 6(f) show that the controller maintains feasibility, and the robot successfully reaches the goal. The visualization in Fig. 6 also shows that the nominal control input was not mapped to the boundary of the feasible set, but to its interior as mentioned in Remark 2. The video is available in supplementary material.
VI-B Case Study 2: AWS Hospital

### v2

@@ V-C2 Chebyshev Ball of a Polytope
The center $c$ and radius $r$ of the largest sphere contained inside a polytope, called Chebyshev ball, can be computed with the following Linear Program (LP)[30, Chapter 8]
@@
where $a_{i}$ and $b_{i}$ are the $i^{th}$ rows of polytope matrices $A,b$.
V-C3 Inscribing Ellipsoid
An ellipoid $E$ is an image of the unit ball under affine transformation $(B,d)$, $E=\{Bu+d~{}|~{}||u||\leq 1\}$ where $B\in S_{++}^{m}$ and $d\in\mathbb{R}^{m\times 1}$. Using the fact that the ellipsoid volume is proportional to $\det B$, the maximum volume ellisoid can be obtained by solving the following convex optimization problem [30, Chapter 8]
$\min_{B,d}\quad$ $\log\det B^{-1}$ (14a)
@@ V-C3 Inscribing Ellipsoid
In practice, Monte Carlo provides most accurate estimates of polytope volume in our simulation studies. The simple MC method in (12) however is known to perform poorly in high-dimensional spaces. Since Monte Carlo estimates are prone to be noisy, this method is also not very suitable for evaluating gradients, which are needed to enforce CBF conditions. While more accurate but relatively expensive MC variations for volume computations[31] are known to exist, we leave the investigation of their suitability for control applications to future work. The implementation with circle approximation is fastest since it only requires solving an LP however it suffers from non-uniqueness of solutions and therefore the LP solver may converge to different solutions for the same state configuration (We comment more on the practical implications of this in Section VI). Moreover, the circle may severely underapproximate the polytope volume. The ellipsoid approximates the volume much more accurately than the circle but is slower to implement. In Section VI, we observe that even with 10 constraints, the SDP can be solved fast enough for real-time implementations.
V-D Smooth Approximation of Polytope Volume
The polytope volume $\mathcal{V}(t,x)$ is not a differentiable function in general. An example configuration that is non-differentiable is shown in Fig. 4. Therefore, $\mathcal{V}(t,x)$ should ideally be formulated as a non-smooth CBF[2] and generalized gradients of the exact volume given by (9) as well as of the ellipse and circle approximations should be used. We defer such an extensive non-smooth analysis to future work. In this work, we utilize gradients of circle and ellipse approximations that are computed using existing tools, details of which are mentioned in Section VI. In the remainder of this section, we propose a smooth underapproximation of (9).
@@ V-D Smooth Approximation of Polytope Volume
Consider the feasible space $\mathcal{U}_{c}(t,x)$ defined in (8) and its volume $\mathcal{V}(t,x)=\textrm{vol}(\mathcal{U}_{c}(t,x))$. Let $H:\mathbb{R}\rightarrow\{0,1\}$ be the Heaveside unit step function such that $H(y)=0$ for $y<0$ and $H(y)=1$ otherwise. Let $H_{s}$ be a differentiable approximation of $H$ such that $H_{s}(z)\leq H(x),\forall z\in\mathbb{R}$. Define $\mathcal{V}_{s}:\mathbb{R}^{+}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ as the following differentiable function
[…]
Then $\mathcal{V}(t,x)\geq\mathcal{V}_{s}(t,x)$ for all $t\in\mathbb{R}^{+},x\in\mathbb{R}^{n}$.
@@
For brevity, we skip the explicit dependency on time in the following. The volume of the polytope (8) is given by
[…]
where $d\mathcal{V}$ is the infinitesimal volume (for example, $d\mathcal{V}=du_{x}du_{y}$ when $m=2$) Note that although $H$ is a discontinuous function, the integral (16b) exists as the points of discontinuity, defined by the boundary of the polytope, form a set of measure 0. Furthermore, since $H(y)\leq H_{s}(y)$ for all $y\in\mathbb{R}$, we have $\mathcal{V}(t,x)\geq\mathcal{V}_{s}(t,x)$. ∎
Note that, although smooth, evaluating the integral (16b) analytically is still non-trivial and requires further numerical techniques such as MC approximation. Due to sensitivity of gradient to the number of samples chosen, we observe that MC, although suitable for obtaining volume, is not suitable for gradient computation. Therefore for lack of space, we only present results with ellipse and circle approximations in the ensuing. Methods to provide stable gradients through (16b) will be explored in future work.
@@ VI Simulation Results
In this section, we present two case studies to evaluate the performance of the proposed volume barrier function. All the simulations are run on a laptop with i9-13905 CPU. The dynamics are propagated using Euler integration with a time step of 0.01s. We use Python 3.11 and JAX’s autodiff feature to compute all required gradients except for the circle and ellipse approximations whose details are mentioned in the next section. Before proceeding with the results, we first mention some important implementation details of our framework.
The sphere and ellipsoid underapproximations of $\mathcal{V}(t,x)$ are computed using LP and SDP respectively, which are convex programs and, therefore, we can exploit sensitivity analysis [32] to compute the gradients of their volume w.r.t the polytope matrices $A,b$. Libraries like cvxpy [33] readily provide an interface to differentiate through convex programs and we use cvxpy’s JAX interface in our simulations. While LP and SDPs are not always differentiable, we do not encounter observable issues in our simulations. We hypothesize that this is because the configurations that are non-differentiable, which are related to instances when a new hyperplane enters or exits the feasible space, are rarely encountered exactly in a discrete-time simulation. A more thorough analysis though is subject to future work. Note that, if any non-differentiable configuration is encountered, cvxpy returns heuristic gradients.
In all our simulations, we first obtain the polytope volume $\mathcal{V}$ and its gradients using one of the proposed methods: smooth approximation or circle/ellipsoid approximation. We then use these values to formulate the CBF conditions and solve the controller (11). Therefore, in the case of circles and ellipsoids, we solve two optimization problems sequentially: first, an LP/SDP, and then the FS-CBF-QP.
VI-A Case Study 1: Valid FS-CBF computation
In this case study, we demonstrate the concept of FS-CBF for navigating in the obstacle environment shown in Fig. 2. A Dubins car model is considered: $\dot{p}_{x}=v\cos\theta,\dot{p}_{y}=v\sin\theta,\dot{\theta}=u$ where $p_{x},p_{y}$ are position, $\theta$ is heading, $v$ is fixed speed, and the angular velocity $u\in[-0.5,0.5]$ is the control input. We consider the constraint functions, also called candidate CBFs, $c_{i}=d_{i}^{2}-d_{min}^{2}$, with $d_{i}$ being the distance to the center of $i^{th}$ obstacle and $d_{min}$ the radius of obstacle. Next, we find valid CBFs $h_{i}$ to each $c_{i}$ using an existing method called refineCBF proposed in [16] that leverages Hamilton-Jacobi (HJ) reachability analysis for designing CBFs. Note that each valid CBF $h_{i}$ does not enforce collision avoidance with an obstacle $j\neq i$. Now we present results for two cases: In Case 1, we start with valid CBFs $h_{i}$ and show their compatible state space in Fig. 5(a). The feasible space volume (of one-dimensional control space in this example) is computed using 100 MC samples using method in Section V-C1. We once again use refineCBF to find a valid CBF on this compatible state space. The resulting FS-CBF is visualized in Fig. 5(b). In Case 2, we start with the candidate CBFs $c_{i}$ and visualize their compatible state space in Fig. 5(c),(e) for two different choices of linear class-$\mathcal{K}$ function parameters. Note that $c_{i}$ are higher-order barriers and their feasible space is defined by (6) for $r_{i}=2$. Next, we find a valid CBF on compatible state space of $c_{i}$ using the refineCBF framework and visualize the resulting FS-CBF in Fig. 5(d),(f). Observe that compatibility is highly sensitive to chosen class-$\mathcal{K}$ parameters. This presents a significant challenge in practice wherein manually tuned candidate CBFs are often employed as finding a valid CBF is not always feasible computationally.
Note that the unsafe region (function value of zero in Fig.5) is larger in Fig.5)(b),(d),(f) compared to Fig.5)(a),(c),(e). This implies that some of the states with non-zero feasible space volume are outside the found forward invariant set and are therefore unsafe. We thus reiterate that CBFs designed independently do not possess any guarantee of compatibility when enforced together.
Figure 5: Compatible space vs refined FS-CBF values at different $X,Y$ and heading $\theta=45^{0}$. The color bar represents the numerical values of FS volume and FS-CBF clipped to the interval [0,1]. (a),(b) Case 1 - initiating with three valid CBFs. (c)-(f) initiating with three candidate CBFs with aggressive (Case 2) and conservative (Case 3) class-$\mathcal{K}$ function parameters for HOCBF. The compatible space (purple) is a subset of the forward invariant set (yellow) as also illustrated in Fig. 3
VI-B Case Study 2: Improving Feasibility and Sensitivity
@@
Figure 6: Case study 1 - humans move towards the right. (a),(b) Paths for the three cases of $k_{v}=1,1.5,2.5$. Blue and red paths in (a) end when CBF-QP becomes infeasible. (c), (d), (e), (f) Feasible space visualization with and without volume barrier function for $k_{v}=1.5$ in (c), (d) and $k_{v}=2.5$ in (e), (f). Note when nominal input is outside the feasible space, the volume barrier function maps it to the interior of feasible space whereas CBF-QP (5) maps it to a point on the boundary.
@@
Consider the scenario shown in Fig. 6 for a robot navigating to its goal location $g$ while avoiding black prohibited zones and performing collision avoidance with humans. This case study serves two purposes. First, we show that the addition of a feasible space volume barrier function enhances compatibility of candidate CBFs. Second, it also helps reduce sensitivity to the hyperparameters of the controller. The robot is modeled as a dynamic unicycle with position $p=[p_{x},p_{y}]$, velocity $v$, heading angle $\psi$ and the following dynamics $\dot{p}_{x}=v\cos\psi,~{}\dot{p}_{y}=v\sin\psi,~{}\dot{v}=a,~{}\dot{\psi}=\omega$. The control inputs are the linear acceleration $a$ and angular velocity $\omega$. The humans move with social force model [34][35], and consider the robot as one of the social agents. We consider the following reference (nominal) controller for the extended unicycle dynamics
@@
where $k_{\omega},k_{p},k_{v}>0$ are user-chosen gains. The three candidate CBFs are designed in similar fashion as in Section VI-A. Fig. 6 shows compares the performance of CBF-QP (5) and FS-CBF-QP (11) for three different values of the gain $k_{v}$. Fig. 6(a)-(b) show the trajectories of the robot, and Figs. 6(c)-6(e) show the variation of polytope volume under three different values of the gain $k_{v}$. We observe that when $k_{v}$ is increased, that is, the nominal controller is more aggressive, the CBF-QP becomes infeasible sooner. FS-CBF-QP on the other hand can maintain feasibility and the robot successfully reaches the goal. as seen in Figs. 6(d), 6(f). The visualization in Fig. 6 also shows that the nominal control input was not mapped to the boundary of the feasible set, but to its interior as mentioned in Remark 2. The video is available on https://www.youtube.com/playlist?list=PLJxod9x5m_V2R5c4MLCCwwe_xliJgzOr2.
The importance of our controller thus lies in enhancing the performance of a given CBF-QP controller. To further corroborate our claim, we perform extensive Monte Carlo simulations by randomizing the robot’s initial state $p_{x},p_{y},v$ and choice of gains $k_{x},k_{v}$. We report the average simulation run times in Table I. The run time is defined as the minimum of 7 seconds (maximum duration of the simulation) and the time at which CBF-QP ($T_{1}$) or FS-CBF-QP ($T_{2}$, reported for different $\alpha_{v}$s) become infeasible. We present results for small as well as large perturbarion in the initial state and gains $k_{x},k_{v}$. In both the cases, we observe that FS-CBF-QP significantly outperforms CBF-QP. Note that infeasibility is not unexpected even for FS-CBF-QP as we only employ candidate CBFs and assume, for sake of implementation, that $\mathcal{V}_{c}=\mathcal{V}$. Since $\mathcal{V}$ is not a valid CBF, our claim is strictly limited to empirical demonstration of enhanced feasibility of QP and reduced sensitivity to user parameters. Designing a CBF for such a complicated scenario with multiple obstacle and heterogenerous agents is an active area of research and no generic solutions exist yet to the best of our knowledge.
$T_{1}$ $T_{2}$ $\alpha_{\mathcal{V}}$ $p_{x}$ $p_{y}$ $v$ $k_{p}$ $k_{v}$
Scene 1 1.67 4.07 0.8 [0.9,1.1] [0.9,1.1] [1.2,1.4] [0.9,1.1] [1.5,1.6]
4.37 1.0
3.53 2.0
Scene 2 1.29 3.04 0.8 [0.9,1.1] [0.9,1.1] [1.2,1.4] [0.9,1.1] [2.5,2.6]
2.13 1.0
1.86 2.0
Scene 3 1.64 4.74 0.8 [0.8,1.2] [0.0,1.5] [1.0,1.5] [0.5,3.0] [0.5,4.0]
4.47 1.0
3.65 2.0
TABLE I: Sensitivity analysis of time to incompatibility for small (Scene 1,2) and large (Scene 3) perturbations of the initial state $p_{x},p_{y},v$ and gains $k_{p},k_{v}$. Variables are sampled with uniform distribution within the specified [lower bound, upper bound].
VI-C Case Study 3: AWS Hospital